The sum of 3 consecutive integers is 60. What is the value of the third integer?
Answer by mananth(16949)

let the integers be x,x+1,x+2
x+(x+1)+(x+2)= 60
3x+3=60
3x=60-3
3x=57
/3
x=19
19,20,21